# Hisham Cholakkal2  Salman Khan1,2,3  Rao Muhammad Anwer2,4

### v1

@@ Abstract
Figure 1: Chronological Wheel of Arabic Poetic Eras. This circular taxonomy visualizes the evolution of Arabic poetry across twelve major historical eras, from the Pre-Islamic and Transitional periods through the Abbasid, Andalusian, and Mamluk dynasties, up to the Modern and Contemporary era. The layout reflects both temporal flow and the rich cultural shifts that shaped poetic expression. Detailed taxonomy by genre, meter, and notable poets presented in Table 2.
@@
Table 1: Comparison of key Arabic NLP datasets. Existing Arabic NLP resources typically address isolated features such as dialectal coverage, QA formats, or classical text processing. In contrast, Fann or Flop uniquely integrates multiple underrepresented dimensions—dialectal diversity, poetic device annotation, verse-level granularity, temporal grounding, and a QA-style evaluation format—positioning it as the first comprehensive benchmark for Arabic poetry understanding. AQMAR Mohit et al. (2012), Tafsir Ahmed et al. (2022), Ashaar Alyafeai et al. (2023), AraBench Sajjad et al. (2020), Arabic-SQuAD Mozannar et al. (2019), ARCD Mozannar et al. (2019), AraBERT Collection Antoun et al. (2020), CAMeL Corpus Abdul-Mageed et al. (2020); Khalifa et al. (2018), Tashkeela Zerrouki and Balla (2017), PADIC Meftouh et al. (2015), MADAR Bouamor et al. (2018).
@@ 1 Introduction
Recent advances in LLMs -such as GPT Chen et al. (2025), LLaMA Touvron et al. (2023), AceGPT Huang et al. (2023), Jais Sengupta et al. (2023), and Falcon Malartic et al. (2024)—have demonstrated impressive multilingual capabilities, including in Arabic. However, most Arabic natural language processing (NLP) benchmarks focus on tasks like sentiment analysis, question answering, or named entity recognition Antoun et al. (2020); Abdul-Mageed et al. (2021); Obeid et al. (2020), typically in MSA or dialectal prose. These benchmarks do not capture the linguistic richness or cultural specificity embedded in Arabic poetry. As LLMs are increasingly evaluated for their ability to handle complex linguistic phenomena, such as metaphor, figurative language, and stylistic nuance, their limitations become evident Liu et al. (2022); Bisk et al. (2020). The FLUTE benchmark Chakrabarty et al. (2022) and the FigLang 2024 workshop FIGLANG202 (2024) have reaffirmed that non-literal language understanding remains a significant challenge. This challenge is particularly acute in Arabic, where poetry is densely layered with intertextuality and cultural symbolism. Arabic poetry thus provides a uniquely demanding testbed for assessing deep linguistic in language models.
To address this gap, we introduce Fann or Flop, the first benchmark dedicated to evaluating LLMs’ understanding of Arabic poetry. Our benchmark includes 6,984 poem-explanation pairs curated from 12 distinct historical poetic eras (see Figure 1), grouped under three major historical periods: pre-Islamic, classical, and modern, and encompasses 21 poetic genres and 31 thematic categories (as shown in Table 1). Each sample is manually verified by native Arabic speakers with domain knowledge to ensure linguistic authenticity and interpretive accuracy. This rich and diverse collection makes Fann a reliable benchmark for evaluating deep cultural and literary reasoning in Arabic NLP. Figure 2 represents the examples from our proposed Fann or Flop dataset, showcasing the diversity of eras, genres, and poetic styles covered.
@@
Table 2: Taxonomy of Arabic Poetic Eras with Genre and Meter Coverage. This table provides a structured overview of twelve major eras in Arabic poetic history, detailing their approximate chronological spans, the most prominent poetic themes (genres) representative of each era, the dominant metrical patterns (Arabic buḥūr) used in poetic composition, and notable poets who exemplify the literary character of their time. The genre column highlights recurring thematic concerns such as satire, elegy, love, nationalism, and religious devotion, while the meter column showcases the classical metrical forms like At-Tawil, Al-Kamel, and Ar-Rojz, along with innovations such as free verse in the modern period. This taxonomy reflects the dynamic interplay between form, content, and historical context in shaping Arabic poetic expression.
Our goal is to provide a diagnostic on how well your language model understands and interprets genuine Classical Arabic. Unlike general text, poetry requires sensitivity to rhetorical devices, metrical patterns, and sociohistorical context—making it a rigorous and culturally grounded indicator of language proficiency. Our benchmark serves as a clear indicator of whether a model has truly been exposed to and internalized high-quality Arabic content, offering a focused lens into its cultural and linguistic depth. We evaluate a range of open-source and commercial LLMs using Fann or Flop and find that, despite strong performance on conventional Arabic tasks, most models struggle with the interpretive depth required by poetry. These findings highlight the need for culturally informed benchmarks that better reflect the depth and diversity of Arabic. We release Fann or Flop as an open-source resource to support the development and evaluation of Arabic-capable language models.
Figure 2: Representative Poetic Samples Across Arabic Literary Eras. This figure presents curated excerpts from Arabic poems spanning key historical eras, illustrating the evolution of language, themes, and stylistic expression. The Pre-Islamic sample reflects tribal valor and rhetorical precision; the Umayyad excerpt captures satire and social commentary; the Abbasid example highlights philosophical introspection and refined metaphorical use; the Transitional era showcases a poetic voice confronting injustice and advocating moral clarity; and the Andalusian selection reveals emotional vulnerability and psychological depth through lyrical expression. Together, these samples provide insight into how Arabic poetry has adapted to diverse historical, cultural, and ideological contexts. Refer to Appendix E.1, Figure 9 for the GPT-4o-generated English translations of the Arabic poetic samples.
@@ 2.1 Dataset Taxonomy
To capture the linguistic, historical, and thematic richness of Arabic poetry, we construct an expert-verified taxonomy that organizes poems across both form and era. As illustrated in Figure 1 and detailed in Table 2, the taxonomy spans twelve distinct poetic eras, ranging from the pre-Islamic period to modern times, and encompasses 21 poetic genres and 31 thematic categories. Each entry reflects the main themes and styles of its time, as well as the historical context, showing how poetic expression changed over the centuries. This structured framework was carefully reviewed and validated by scholars specializing in Arabic language and literature to ensure both linguistic accuracy and contextual relevance. Their expertise helped align the taxonomy with established literary traditions while accommodating the nuances of classical and modern poetic forms. Beyond its utility for literary and philological analysis, the taxonomy serves as a robust foundation for computational modeling. It enables more precise automatic genre classification, facilitates temporal contextualization across different eras of Arabic poetry, and supports culturally informed natural language processing (NLP) research. By embedding cultural and historical dimensions, our dataset enhances the interpretability and relevance of Arabic NLP research.
@@ 2.2 Data Collection
We curated Arabic poems from a well-established digital archive [^3] https://arabic-poetry.net, which hosts a broad spectrum of poets, genres, and historical periods. A custom web scraper was developed to extract the poem texts along with associated metadata, including poet name, historical era, and poem title. The resulting dataset spans 12 distinct eras, from the Pre-Islamic period to the Modern era, and reflects a diverse range of poetic forms and styles. All entries were curated in accordance with our expert-verified taxonomy (see Table 2), ensuring consistency across genre and era classifications. This structured approach preserves both the linguistic richness and historical specificity of Arabic poetry, yielding a valuable resource for research in both Arabic NLP and digital literary studies. By aligning each sample with a well-defined literary context, the dataset enables more accurate model evaluation and supports culturally grounded language understanding.
@@ 2.3 Data Filtering and Verification
    Expert Validation of Labels: All genre and era annotations were reviewed by Arabic language and literature experts. This validation step ensured that each poem was accurately categorized in accordance with the taxonomy introduced in Section 2.
@@ 3 Fann or Flop Benchmark Evaluation
Table 4 shows era-wise performance of closed and open-source models on Arabic poem understanding using BERTScore, which captures semantic similarity with human explanations. Closed models like GPT-4o and Gemini variants perform consistently well, especially on modern and recent historical eras. In contrast, open models such as Deepseek-V3 and Aya-Expanse perform reasonably on some eras but struggle with older poetic forms like Pre-Islamic and Umayyad due to their complex language and cultural depth. This highlights that while current models are effective on Modern Standard Arabic, they face challenges with classical Arabic. A complementary analysis using Textual Entailment is included in the Appendix (refer Table 17), further supporting these findings.
@@ 4 Conclusion
Arabic poetry represents one of the richest and most culturally nuanced forms of expression within the Arabic language, characterized by layered meanings, stylistic diversity, and deep historical roots. In this paper, we introduced Fann or Flop, the first benchmark specifically developed to evaluate the capabilities of large language models (LLMs) in understanding Arabic poetry across twelve historical eras, spanning from pre-Islamic to contemporary periods, and encompassing a broad spectrum of poetic genres and metrical forms. Our benchmark includes carefully curated diagnostic questions aimed at assessing semantic comprehension, metaphorical interpretation, prosodic awareness, and sensitivity to cultural contexts. Through extensive evaluation, we demonstrated that despite strong performances on standard Arabic language tasks, state-of-the-art LLMs consistently struggle with the interpretative and culturally embedded dimensions of Arabic poetic texts. By releasing Fann or Flop as an open-source resource, we aim to catalyze further research, promote rigorous assessment methodologies, and drive advancements in linguistically and culturally sophisticated Arabic language modeling.
@@ 5 Limitations and Societal Impact
While this research demonstrates the potential of LMMs in enhancing historical documentation, the quality of generated descriptions depends on the completeness and accuracy of the input data. In cases where historical records are fragmented or ambiguous, AI-generated text may lack full contextual depth. Additionally, biases present in training data can influence how models interpret and describe cultural artifacts, necessitating continuous evaluation and expert validation to ensure historical accuracy and cultural sensitivity. Despite these challenges, this research contributes to cultural heritage preservation, educational accessibility, and AI-driven humanities research. By digitizing and enriching historical records, it enables wider public engagement with history, supports museum digitization efforts, and provides a foundation for future advancements in AI-assisted historical analysis, bridging the gap between technology and human expertise in understanding our collective past.
@@ Appendix A Appendix
This appendix provides supplementary material to support our study of Arabic poetry understanding in language models. It includes three key sections: (1) a brief review of related work in Arabic NLP, highlighting progress in benchmark development and the gaps our work addresses; (2) detailed data statistics, including token distribution, genre coverage, and temporal representation across poetic eras; and (3) a selection of qualitative examples from the Fann or Flop benchmark, illustrating its richness and the interpretive challenges they pose. Together, these components reinforce the linguistic, historical, and cultural depth of our dataset and benchmark evaluation.
@@ Appendix C Fann or Flop Data Statistics
Tables 5 to 7 provide a breakdown of the number of poems attributed to prominent Arabic poets across different historical eras. Each table is dedicated to one era:
Table 5: Pre-Islamic era; Table 6 Transitional (Early Islamic) period; Table 8: Islamic era; Table 9: Umayyad era; Table 10: Abbasid era; Table 12: Fatimid Dynasty; Table 13: Andalusian era; Table 15: Mamluk Dynasty; Table 11: Between 2 Dynasties; Table 16: Ottoman era; Table 7: Modern era.
@@ D.1 Model Generation Prompt
Figure 8: System prompt used for LLM-Judge evaluation of verse-by-verse poem explanations. LLM OpenAI (2024) compare AI-generated outputs with original explanations and assign overall scores for faithfulness to meaning and fluency in Modern Standard Arabic, following clearly defined criteria. The structured format ensures consistency and reliability across evaluations.
@@ Appendix E Additional Examples: Qualitative, Translated, and Quantitative Insights
In the following section, we present a more detailed evaluation of the Textual-Entailment (Refer Table 17) metric across the twelve historical eras, comparing both open-source and closed-source models on this dimension. To support comprehensive engagement and a clearer understanding of the data evaluated, we also include selected English translations of Arabic poetic samples as well as additional qualitative Arabic examples. These additions offer deeper insight into the linguistic diversity, thematic range, and overall quality of the dataset used in our analysis.

### v2

# Hisham Cholakkal1  Salman Khan1,2  Rao Muhammad Anwer1,3

@@ Abstract
Figure 1: Chronological Wheel of Arabic Poetic Eras. This circular taxonomy visualizes the evolution of Arabic poetry across 12 major historical eras, from the Pre-Islamic and Transitional periods through the Abbasid, Andalusian, and Mamluk dynasties, up to the Modern era. The layout reflects both temporal flow and the rich cultural shifts that shaped poetic expression. Detailed taxonomy by genre, meter, and notable poets presented in Table 2.
@@
Table 1: Comparison of key Arabic NLP datasets. Existing Arabic NLP resources typically address isolated features such as dialectal coverage, QA formats, or classical text processing. In contrast, Fann or Flop uniquely integrates multiple underrepresented dimensions (i.e. dialectal diversity, poetic device annotation, verse-level granularity, temporal grounding, and a QA-style evaluation format) positioning it as the first comprehensive benchmark for Arabic poetry understanding. AQMAR Mohit et al. (2012), Tafsir Ahmed et al. (2022), Ashaar Alyafeai et al. (2023), AraBench Sajjad et al. (2020), Arabic-SQuAD Mozannar et al. (2019), ARCD Mozannar et al. (2019), AraBERT Collection Antoun et al. (2020), CAMeL Corpus Abdul-Mageed et al. (2020); Khalifa et al. (2018), Tashkeela Zerrouki and Balla (2017), PADIC Meftouh et al. (2015), MADAR Bouamor et al. (2018).
@@ 1 Introduction
Recent advances in LLM, such as GPT Chen et al. (2025), LLaMA Touvron et al. (2023), AceGPT Huang et al. (2023), Jais Sengupta et al. (2023), and Falcon Malartic et al. (2024), have demonstrated impressive multilingual capabilities, including Arabic. However, most Arabic natural language processing (NLP) benchmarks focus on tasks such as sentiment analysis, question answering, or recognition of named entities Antoun et al. (2020); Abdul-Mageed et al. (2021); Obeid et al. (2020), typically in MSA or dialectal prose. These benchmarks often miss the linguistic depth and cultural nuances that are inherent in Arabic poetry. As LLMs are increasingly evaluated for their ability to handle complex linguistic phenomena, such as metaphor, figurative language, and stylistic nuance, their limitations become evident Liu et al. (2022); Bisk et al. (2020). The FLUTE benchmark Chakrabarty et al. (2022) and the FigLang 2024 workshop FIGLANG202 (2024) have reaffirmed that non-literal language understanding remains a significant challenge. This challenge is particularly acute in Arabic, where poetry is densely layered with intertextuality and cultural symbolism. Arabic poetry thus provides a uniquely demanding testbed for assessing deep linguistic in language models.
To address this gap, we introduce Fann or Flop, the first benchmark dedicated to evaluating LLMs’ understanding of Arabic poetry. Our benchmark comprises 6,984 poem-explanation pairs curated from 12 distinct historical poetic eras (see Figure 1), which can be broadly seen as spanning three major historical periods: pre-Islamic, classical, and contemporary. It covers 14 poetic genres and includes a range of metrical forms, as detailed in Table 1. Each sample is manually verified by native Arabic speakers with domain knowledge to ensure linguistic authenticity and interpretive accuracy. This rich and diverse collection makes Fann or Flop a reliable benchmark for evaluating deep cultural and literary reasoning in Arabic NLP. Figure 2 represents the examples from our proposed Fann or Flop dataset, showcasing the diversity of eras, genres, and poetic styles covered.
@@
Table 2: Taxonomy of Arabic Poetic Eras with Genre and Meter Coverage. This table provides a structured overview of 12 major eras in Arabic poetic history, detailing their approximate chronological spans, the most prominent poetic themes (genres) representative of each era, the dominant metrical patterns (Arabic buḥūr) used in poetic composition, and notable poets who exemplify the literary character of their time. The genre column highlights recurring thematic concerns such as satire, elegy, love, nationalism, and religious devotion, while the meter column showcases the classical metrical forms like At-Tawil, Al-Kamel, and Ar-Rojz, along with innovations such as free verse in the modern period. This taxonomy reflects the dynamic interplay between form, content, and historical context in shaping Arabic poetic expression.
Our goal is to provide a diagnostic on how well your language model understands and interprets genuine Classical Arabic. Unlike general text, poetry requires sensitivity to rhetorical devices, metrical patterns, and sociohistorical context, making it a rigorous and culturally grounded indicator of language proficiency. Our benchmark serves as a clear indicator of whether a model has truly been exposed to and internalized high-quality Arabic content, offering a focused lens into its cultural and linguistic depth. We evaluate a range of open-source and commercial LLMs using Fann or Flop and find that, despite strong performance on conventional Arabic tasks, most models struggle with the interpretive depth required by poetry. These findings highlight the need for culturally informed benchmarks that better reflect the depth and diversity of Arabic. We release Fann or Flop as an open-source resource to support the development and evaluation of Arabic-capable language models.
Figure 2: Representative Poetic Samples Across Arabic Literary Eras. This figure presents curated excerpts from Arabic poems spanning key historical eras, illustrating the evolution of language, themes, and stylistic expression. The Pre-Islamic sample reflects tribal valor and rhetorical precision; the Umayyad excerpt captures satire and social commentary; the Abbasid example highlights philosophical reflection and refined metaphorical use; the Transitional era showcases a poetic voice confronting injustice and advocating moral clarity; and the Andalusian selection reveals emotional openness and psychological depth through lyrical expression. Together, these samples provide insight into how Arabic poetry has adapted to diverse historical, cultural, and ideological contexts. Refer to Appendix E.1, Figure 9 for the GPT-4o-generated English translations of the Arabic poetic samples.
@@ 2.1 Dataset Taxonomy
To capture the linguistic, historical, and thematic richness of Arabic poetry, we construct an expert-verified taxonomy that organizes poems across both form and era. As illustrated in Figure 1 and detailed in Table 2, the taxonomy traces 12 distinct poetic eras, from the pre-Islamic period to modern times, encompassing 14 genres that capture the dominant styles, concerns, and historical contexts of each era. It illustrates how poetic expression evolved over the centuries.
This structured framework was carefully reviewed and validated by scholars specializing in Arabic language and literature to ensure both linguistic accuracy and contextual relevance. Their expertise helped align the taxonomy with established literary traditions while accommodating the nuances of classical and modern poetic forms. Beyond its utility for literary and philological analysis, the taxonomy serves as a robust foundation for computational modeling. It enables more precise automatic genre classification and facilitates temporal contextualization across different eras of Arabic poetry, thereby supporting culturally informed and interpretable Arabic NLP research.
@@ 2.2 Data Collection
We curated Arabic poems from a well-established digital archive [^3] https://arabic-poetry.net, which hosts a broad spectrum of poets, genres, and historical periods. A custom web scraper was developed to extract the poem texts along with associated metadata, including the poet’s name, historical era, genre, meter, and poem’s title. The resulting dataset extends across 12 distinct eras, from the pre-Islamic period to the modern era, and reflects a diverse range of poetic forms and styles. All entries were curated following our expert-verified taxonomy (see Table 2), ensuring consistency across genre and era classifications. This structured approach preserves both the linguistic richness and historical specificity of Arabic poetry, providing a valuable resource for research in both Arabic NLP and digital literary studies. By aligning each sample with a well-defined literary context, the dataset enables more accurate model evaluation and supports culturally grounded language understanding.
@@ 2.3 Data Filtering and Verification
    Expert Validation of Labels: All genre and era annotations were reviewed by Arabic language and literature experts. This validation step ensured that each poem was accurately categorized in accordance with the taxonomy introduced in Section 2.1.
@@ 3 Fann or Flop Benchmark Evaluation
Table 4 shows era-wise performance of closed and open-source models on Arabic poem understanding using BERTScore, which captures semantic similarity with human explanations. Closed models like GPT-4o and Gemini variants perform consistently well, especially on modern and recent historical eras. In contrast, open models such as Deepseek-V3 and Aya-Expanse perform reasonably on some eras but struggle with older poetic forms like Pre-Islamic and Umayyad due to their complex language and cultural depth. This highlights that while current models are effective on MSA, they face challenges with classical Arabic. A complementary analysis using Textual Entailment is included in the Appendix (refer Table 17), further supporting these findings.
@@ 4 Conclusion
Arabic poetry represents one of the richest and most culturally nuanced forms of expression within the Arabic language, characterized by layered meanings, stylistic diversity, and deep historical roots. In this paper, we introduced Fann or Flop, the first benchmark specifically developed to evaluate the capabilities of LLMs in understanding Arabic poetry across 12 historical eras, spanning from pre-Islamic to contemporary periods, and encompassing a broad spectrum of poetic genres and metrical forms. Our benchmark includes carefully curated diagnostic questions aimed at assessing semantic comprehension, metaphorical interpretation, prosodic awareness, and sensitivity to cultural contexts. Through extensive evaluation, we demonstrated that despite strong performances on standard Arabic language tasks, state-of-the-art LLMs consistently struggle with the interpretative and culturally embedded dimensions of Arabic poetic texts. By releasing Fann or Flop as an open-source resource, we aim to encourage further research, promote rigorous assessment methodologies, and support advancements in linguistically and culturally rich Arabic language modeling.
@@ 5 Limitations and Societal Impact
While Fann or Flop provides a rigorous framework for evaluating LLMs’ understanding of Arabic poetry, it has several limitations. The benchmark covers only a portion of the broader Arabic poetic tradition, as some poems could not be included due to missing metadata, unclear authorship, or lack of reliable era or genre annotations. Additionally, poetry often invites multiple valid interpretations, which current evaluation metrics may not fully capture, even with expert-curated references. Expanding the dataset to include more diverse annotations, as well as dialectal and regional poetic forms, remains a key area for future work.
On the societal front, this benchmark contributes to the preservation and computational accessibility of Arabic literary heritage by positioning poetry as a meaningful testbed for language understanding. By promoting the development of culturally informed and linguistically grounded models, Fann or Flop encourages more inclusive and context-sensitive NLP. Nonetheless, as with any system trained on culturally rich and potentially sensitive material, there is a risk of misinterpretation or misuse. Ensuring transparency, human oversight, and responsible deployment is essential to safeguard the ethical impact of this work, especially in educational, literary, and public-facing applications.
@@ Appendix A Appendix
This appendix provides supplementary material to support our study of Arabic poetry understanding in language models. It includes four key sections: (1) a brief overview of related work in Arabic NLP, highlighting recent progress in benchmark development and the specific gaps our work addresses; (2) detailed dataset statistics, including token distribution, genre coverage, and temporal representation across poetic eras; (3) additional details on the prompts used for model generation and evaluation; and (4) a selection of qualitative examples from the Fann or Flop benchmark that illustrate its richness and the interpretive challenges it presents. Together, these components underscore the linguistic, historical, and cultural depth of our dataset and evaluation framework.
@@ Appendix C Fann or Flop Data Statistics
Tables 5 to 16 provide a breakdown of the number of poems attributed to prominent Arabic poets across different historical eras. Each table is dedicated to one era:
Table 5: Pre-Islamic era; Table 6 Transitional (Early Islamic) period; Table 7: Modern era; Table 8: Islamic era; Table 9: Umayyad era; Table 10: Abbasid era; Table 11: Between 2 Dynasties; Table 12: Fatimid Dynasty; Table 13: Andalusian era; Table 14: Ayybid era. Table 15: Mamluk Dynasty; Table 16: Ottoman era.
@@ D.1 Model Generation Prompt
Figure 8: System prompt used for LLM-Judge evaluation of verse-by-verse poem explanations. LLM OpenAI (2024) compare AI-generated outputs with original explanations and assign overall scores for faithfulness to meaning and fluency in MSA, following clearly defined criteria. The structured format ensures consistency and reliability across evaluations.
@@ Appendix E Additional Examples: Qualitative, Translated, and Quantitative Insights
In the following section, we present a more detailed evaluation of the Textual-Entailment (Refer Table 17) metric across the 12 historical eras, comparing both open-source and closed-source models on this dimension. To support comprehensive engagement and a clearer understanding of the data evaluated, we also include selected English translations of Arabic poetic samples as well as additional qualitative Arabic examples. These additions offer deeper insight into the linguistic diversity, thematic range, and overall quality of the dataset used in our analysis.